# Implicit In-context Learning

### v1

@@ 4.4 Ablation Study
In this section, we conduct a comprehensive ablation study on the module, layer, and token position of injection, as well as the noise scale. We also explore various vectorization and injection formulas to highlight the rationale behind our designs.
Target Module I2CL extracts and injects activations at both the MHA and MLP modules. To identify the contribution of each module, we test using either MHA or MLP independently and also consider leveraging the hidden state at each layer. As shown in Table 8, extracting and injecting context vectors at either MHA or MLP proves beneficial, with MLP showing a clear advantage. We hypothesize this is due to the additional engagement of information stored in the MLP weights. However, targeting the hidden state does not lead to improvement, likely due to the accumulation effect that complicates harnessing the scales of hidden states.
Target Layer I2CL encompasses all layers in a large language model (LLM), eliminating the need for a layer selection process, which is often a designated hyperparameter. This model-agnostic approach not only simplifies the setup but also enhances performances, as evidenced in Table 8. We divide the model into three sub-parts—early, middle, and late—each containing one-third of the total layers. We then apply I2CL only within the target layer range. The middle and late layers prove more effective than the early layers, and injecting across all layers provides a clear performance boost, highlighting the importance of fusing information at all levels of abstraction.
@@
Noise Scale Here, we evaluate the impact of noise strength during the calibration phase. As demonstrated in Table 8, an appropriate noise scale leads to a clear performance gain. Conversely, an improper noise scale will disrupt information propagation during inference, resulting in deteriorated outcomes. Through empirical testing, we identify $\tau=0.001$ as a suitable scale and have applied it across all tasks.
Vectorization Method For context vectorization, we treat each demonstration example separately and post-merge them using an element-wise algorithmic average. A straightforward alternative is to concatenate all demonstration examples as in typical ICL, and extract the context vector from the end residual stream. Furthermore, we evaluate another post-fusion strategy that applies PCA to extract the first principal component vector as the context vector. According to Table 8, the concatenation method significantly underperforms the post-fusion counterparts. The PCA variant achieves a negligible improvement; therefore, we employ the average operator by default for its simplicity.
@@
Injection Formula I2CL utilizes a linear combination to blend context vectors with query activations. Let ${\bm{v}}$ denote the context vector and ${\bm{a}}$ indicates activation; we use $\lambda{\bm{c}}+\beta{\bm{a}}$. One simplification is to view the injection process as solely adding the context vector to the activation: $\lambda{\bm{v}}+{\bm{a}}$ with $\lambda>0$. Another common formula involves constraining the sum of linear coefficients to one and allowing a separate scale factor: $(\lambda{\bm{v}}+(1-\lambda){\bm{a}})\times\beta,\beta>0$. According to Table 8, a linear combination with no constraints achieves the best overall performance. Therefore, we conjecture that it is critical to allow not only information addition but also deletion, i.e., permitting a negative sign for the linear coefficient, and to scale each vector independently. Observations in Figure 4 (right) support our conjecture.
5 Discussion
A Hypothetical Two-Stage Interpretation of ICL  The effectiveness of I2CL suggests a conceptual two-stage internal mechanism within ICL. In the first stage, the information from each demonstration example is encapsulated within its respective end residual stream. In the second stage, this encapsulated information is accessed and integrated into the residual streams of the query. Commonly, both stages are seamlessly executed by non-linear multi-head attention modules within transformer layers. I2CL reveals that the end residual stream of a sequence functions as an effective anchor to condense context information, and by leveraging activation vectors from it, we can effectively approximate an attention-based information fusion scheme using only linear operations. We hypothesize that this phenomenon arises due to the emergence of certain linear representations during the pre-training stage. These linear properties enable a vector interface that facilitates the formation and transportation of functions through simple linear operations. While we do not claim the above interpretation as a rigorous internal mechanism of ICL, it aligns with our empirical observations and we hope it can inspire future work aimed at a deeper understanding of the ICL.
Potential Design Although the end residual stream contains a rich repository of information for the given token sequence, other token positions may also possess valuable attributes, as noted in recent studies [Hendel et al., 2023, Todd et al., 2024, Liu et al., 2024]. Recent research also highlights the importance of formatting tokens during information propagation [Wang et al., 2023, Bai et al., 2024]. We believe our approach could benefit from a more sophisticated design of demonstration vectorization. Additionally, while we currently use a single scalar to gauge the strength of aggregated multi-head attention, allowing finer granularity—such as separate strength scalars for different attention heads—might enhance our system’s performance, albeit at the cost of increased parameters.
@@ 5 Discussion
Limitations and Future Work Our research is subject to several limitations. Firstly, we confine the scope of our initial exploration to standard classification tasks, leaving more sophisticated tasks for future research. It would be intriguing to extend I2CL to the realm of open-ended generation tasks and those involving multi-hop reasoning processes. Secondly, I2CL necessitates access to and caching of intermediate activations from language models, which may not be feasible with state-of-the-art commercial models (e.g., GPT-4, Gemini, Claude3). Lastly, we have evaluated I2CL on relatively modest-sized models, and scaling our evaluations to models with hundreds of billions of parameters could yield additional insights.
@@ 6 Conclusion
In this study, we introduce Implicit In-context Learning (I2CL), a simple and novel framework that integrates a minimal set of demonstration examples within the activation space of LLMs. Diverge from ICL, I2CL eliminates the need for caching the latents of demonstration tokens and replaces the non-linear information fusion process with linear operations. Therefore, I2CL reduces both computational and memory expenses during inference to that of zero-shot level. Moreover, I2CL is validated to be robust against token space variations, and it facilitates a novel representation of “task-ids”, enhancing task similarity detection and enabling effective transfer learning. Empirical evidence on nine real-world tasks across three different models demonstrates I2CL as a more efficient and robust alternative to ICL. Through a set of in-depth analyses, we shed light on the internal mechanics of I2CL and provide further insights into the working scheme of ICL.
@@ Acknowledgments
We extend our sincere gratitude to Prof. Diyi Yang for her constructive suggestions on the empirical design of this study. Her expertise and insights were invaluable to this research.
Appendix A Templates
@@ Appendix B Reproduction Details
The code is publicly available at: https://github.com/LzVv123456/I2CL
B.1 Implementation of Comparable Methods.
Noise Vector We replace context vectors with random noise vectors while keeping all other settings identical to I2CL.
Soft Prompt For the implementation of soft-prompt method [Lester et al., 2021], we allow one extra learnable token ($P=1$) per layer, and apply soft prompt across all layers. We also attempt to use more learnable tokens, resulting in poorer performance due to overfitting. For optimization, we conduct a simple grid search on SST-2 to determine an optimal learning rate of $0.1$, which we then apply across all tasks. All other configurations remain as specified in the experimental section.
@@
Task Vector We replicate the task vector method originally studied by Hendel et al. [2023], which was evaluated using a set of toy datasets, within our experimental framework. To generate the task vector, we append a random extra query after the concatenated demonstration examples (five per class) to simulate the dummy query they utilized. We then extract the hidden state from the token position responsible for the next token prediction to serve as the task vector. A hold-out dataset with $32$ extra examples is used to select the best layer for extraction and replacement for each task, following the original practice.
B.2 Implementation of Transfer Learning
Algorithm 1 details the transfer learning method of I2CL. In implementation, we set $h=0.8$ and $\tau=0.5$.
@@ B.2 Implementation of Transfer Learning
Appendix C Additional Experiments
C.1 Results of Other Architectures
Here, we show results of I2CL under architecture GPT-J and GPT2-XL in Table 10 and Table 11, respectively.
Table 10: Comparison between I2CL and other methods on GPT-J-6B.
@@ C.1 Results of Other Architectures
Table 11: Comparison between I2CL and other comparable methods on GPT2-XL. AGnews and DBPedia are not evaluated due to the limitation of GPT2-XL’s context window size.
@@
C.2 Trend of I2CL Under Varying Shots
In our standard empirical setup, we select five examples per class to create the demonstration group. This approach prompts an important question about how I2CL behaves under different shot configurations. To investigate this, we varied the number of shots per class from $1$ to $10$ and compared the performance of I2CL with that of ICL. As illustrated in Figure 5, I2CL and ICL perform similarly when the number of demonstration examples is limited. However, while the performance of ICL plateaued at approximately five shots per class, our model continues to show the capacity to benefit from an increased number of demonstration examples.
Figure 5: Data scaling trend of ICL and I2CL on Llama2-7B.
Appendix D Additional Visualization
@@
Additional Plots of Calibrated Linear Coefficients Here, we present the calibrated coefficients for all tasks we used, illustrating variations and patterns across different configurations.
Figure 6: Calibrated coefficients for SST-2.
Figure 7: Calibrated coefficients for SST-5.
Figure 8: Calibrated coefficients for TREC.
Figure 9: Calibrated coefficients for AGNews.
Figure 10: Calibrated coefficients for Subj.
Figure 11: Calibrated coefficients for MR.
Figure 12: Calibrated coefficients for DBPedia.
Figure 13: Calibrated coefficients for EmoC.
Figure 14: Calibrated coefficients for HateSpeech18.

### v2

@@ 3.4 Ablation Study
In this section, we conduct a comprehensive ablation study on the module, layer, and token position of injection, as well as the noise scale. We also explore various vectorization and injection formulas to highlight the rationale of our designs.
Target Module I2CL extracts and injects activations at both MHA and MLP modules. To identify the contribution of each module, we test using either MHA or MLP independently, and we also consider leveraging the hidden state at each layer. As shown in Table 8, extracting and injecting context vectors at either MHA or MLP proves beneficial, with MLP showing a clear advantage. We hypothesize this is due to the additional engagement of information stored in the MLP weights. However, targeting the hidden state does not lead to any improvement, likely due to the accumulation effect which complicates the optimization process within self-calibration stage.
Target Layer I2CL encompasses all layers in a large language model (LLM), eliminating the need for specifying layer indexes. This model-agnostic approach not only simplifies the setup but also enhances performance, as evidenced in Table 8. We divide the model into three sub-parts—early, middle, and late—each containing one-third of the total layers. We then apply I2CL only within the target layer range. The middle and late layers prove more effective than the early layers, and injecting across all layers provides a clear performance boost, highlighting the importance of fusing information at all levels of abstraction.
@@
Noise Scale We evaluate the impact of noise strength during the calibration phase. As demonstrated in Table 8, an appropriate noise scale leads to a clear performance gain. Conversely, an improper strength can disrupt the propagation of information at inference, resulting in deteriorated outcomes. We empirically identify $\tau=0.001$ as a suitable scale and have applied it across all tasks and architectures.
@@
Injection Formula I2CL utilizes a linear combination to blend context vectors with query activations. Let ${\bm{v}}$ denote the context vector and ${\bm{a}}$ indicates activation; we use $\lambda{\bm{c}}+\beta{\bm{a}}$. One simplification is to view the injection process as solely adding the context vector to the activation: $\lambda{\bm{v}}+{\bm{a}}$ with $\lambda>0$ as done in Liu et al. (2024). Another common formula involves constraining the sum of linear coefficients to one and allowing a separate scale factor: $(\lambda{\bm{v}}+(1-\lambda){\bm{a}})\times\beta,\beta>0$. According to Table 8, a linear combination with no constraints achieves the best overall performance. Therefore, we conjecture that it is critical to allow not only information addition but also deletion, i.e., permitting a negative sign for the linear coefficient, and to scale each vector independently. Observations in Figure 5 (right) corroborate this conjecture.
4 Related Work
Understanding In-context Learning Besides enhancing ICL, the exploration of ICL’s internal mechanisms has attracted significant research attention. akyürek2023learning draw parallels between ICL and gradient descent in linear regression tasks, suggesting a fundamental alignment with classical optimization methods. Complementary perspectives from Von Oswald et al. (2023) and Dai et al. (2023) conceptualize ICL as a form of meta-optimization, further enriching our understanding of its operational basis. Concurrently, Xie et al. (2022) interpret ICL through the lens of implicit Bayesian inference, proposing a probabilistic foundation for the learning process. Wei et al. (2023) and Olsson et al. (2022) respectively attribute ICL’s capabilities to the establishment of input-label correspondences and the identification of so-called induction heads, highlighting the intricate interplay between data representation and model interpretability. Most recently, Label-as-Anchors (Wang et al., 2023) inspects ICL from an information flow perspective and leverages anchor tokens to perform token reduction. Similarly, I2CL also enhances the efficiency of ICL, but distinguishes itself by circumventing the need for caching latents of anchor tokens or employing multi-head attention, thereby reducing the inference cost to that of the zero-shot.
Activation/representation Engineering An emerging research field, termed activation/representation engineering, closely relates to our study. Recent endeavors by Merullo et al. (2023); Turner et al. (2023); Zou et al. (2023) have unveiled the phenomenon of steering vectors, which can be derived from a positive and negative text pair and used to steer the content generation of LLMs. Steering vectors can also be learned via gradient descent (Subramani et al., 2022). Liu et al. (2024) applies these insights to bolster LLM safety while Li et al. (2023) explores their utility in eliciting more truthful responses from LLMs. Steering vector has also been adapted to multi-modal domain to reduce hallucination (Li et al., 2025). To better understand the inner mechanism of activation engineering, the linear representation hypothesis has been studied and discussed in Li et al. (2021); Hernandez et al. (2024); Park et al. (2023). Central to this discourse, the idea of task/function vector (Hendel et al., 2023; Todd et al., 2024) resonates with the core premise of I2CL. Both methodologies extract task/function vectors from the demonstration examples and use them to improve zero-shot performance. I2CL stands out not only due to its superior performance, but also through its simplicity, namely, its avoidance of paired demonstrations, and the need for task- or architecture-specific hyperparameter tuning, such as selecting attention heads through causal mediation or determining target layers via extra validation sets.
Prompt Compression for LLMs Prompt compression (Chang et al., 2024) is also related to our work, as it shares a similar goal of improving the inference efficiency of LLMs. However, prompt compression methods emphasize on reducing the length of excessive contexts by learning an additional amortized compressor (Ge et al., 2023; Chevalier et al., 2023; Yen et al., 2024), while I2CL transforms the standard few-shot learning approach into a zero-shot manner at test time, treating the LLM itself as a context vector generator. Most prompt compression methods operate by compressing a long prompt sequence into a set of compact soft prompts, leveraging the attention mechanism to further propagate context information. In contrast, I2CL generates a context vector from the activation space and intervenes directly in the residual streams.
5 Limitations
@@
I2CL is subject to several limitations. First, we confine the scope of this initial exploration to standard text classification tasks, leaving more sophisticated tasks for future research. It is non-trivial to further extend I2CL to the realm of open-ended generation tasks and those involving multi-hop reasoning processes. Second, I2CL necessitates access to and caching of intermediate activations from language models, which may not be feasible with state-of-the-art commercial models (e.g., GPT-4, Gemini, Claude3). Thirdly, limited by the computational resources, we evaluated I2CL on modest-sized models, and further scaling the evaluation to commercial size models could yield additional insights.
@@ 6 Conclusion
In this study, we introduce Implicit In-context Learning (I2CL), a simple and novel framework that integrates a minimal set of demonstration examples within the activation space of LLMs. Diverging from ICL, I2CL eliminates the need for caching the latents of demonstration tokens and replaces the non-linear information fusion process (i.e., attention head) with linear operations. Therefore, I2CL reduces both computational and memory expenses during inference to that of zero-shot level. Moreover, I2CL is validated to be robust against token space variations, and it facilitates a novel representation of task-ids which enhances task similarity detection and fosters transfer learning. Empirical evidence on nine real-world tasks across three different models suggests the potential of I2CL as a more efficient and robust alternative to ICL for constrained scenarios. Through a set of in-depth analyses and ablations, we also shed light on the internal working mechanics of I2CL.
@@ Acknowledgments
We extend our sincere gratitude to Prof. Diyi Yang for her constructive suggestions on the motivation and empirical design of this study. Her expertise and insights were invaluable to this research.
Appendix A Prompting Templates
@@ Appendix B Reproduction Details
B.1 Implementation of Baseline Methods.
Noise Vector. In this baseline method, We simply replace context vectors with random noises while keeping all other settings identical to I2CL.
@@
Task Vector (Hendel et al., 2023). We replicate the task vector method which was initially evaluated on a set of toy datasets. To generate the task vector, we append a random extra query after the concatenated demonstration examples (five per class) to simulate the dummy query they utilized. We then extract the hidden state from the token position responsible for the next token prediction to serve as the task vector. A hold-out dataset with $32$ extra examples is used to select the best layer for extraction and replacement for each task, following the original method.
Table 10: Strength scalar $\lambda$ for ICV.
Task $\lambda$ SST-2 0.01 SST-5 0.02 TREC 0.02 AGNews 0.0001 Subj 0.02 HateSpeech18 0.02 DBPedia 0.001 EmoC 0.02 MR 0.02
In-contex Vector (Liu et al., 2024). ICV method is primarily crafted for open-end generation tasks, and it leverages positive and negative demonstration pairs to generate the steering vector. In our scenario, we set sentence (or question) in a demonstration example as negative and its corresponding answer as positive counterpart to extract in-context vector. We then manually search the strength scalar of injection for each task. Specifically, we first search in a log scale covering $\lambda\in[0.0001,0.001,0.01,0.1,1.0]$, followed by a more fine-grained search between $0.01$ and $0.1$. The best strength scalars we used are reported in Table 10. We refer readers to https://github.com/shengliu66/ICV for more technique details.
AutoCompressors (Chevalier et al., 2023). We directly test the officially released model from here: https://github.com/princeton-nlp/AutoCompressors on our tasks. All hyper-parameters are set by default values. Evaluation protocols remain the same as ours.
ICAE (Ge et al., 2023). We directly test the officially released model from here: https://github.com/getao/icae on our tasks. All hyper-parameters are set by default values. Evaluation protocols remain the same as ours.
CEPE (Yen et al., 2024). We directly test the officially released model from here: https://github.com/princeton-nlp/CEPE on our tasks. All hyper-parameters are set by default values. Evaluation protocols remain the same as ours.
Prompt-tuning (Lester et al., 2021). For the implementation of prompt-tuning method, we allow one extra learnable token ($P=1$) per layer, and apply learnable prompts across all layers. We also attempt to use more learnable tokens, resulting in poorer performance due to overfitting. For optimization, we conduct a simple grid search on SST-2 to determine an optimal learning rate of $0.1$, which we then apply across all tasks. All other configurations remain as specified in the experimental section.
LoRA (Hu et al., 2021) and IA3 (Liu et al., 2022). We use implementations from HuggingFace PEFT library for both PEFT methods. Learning rate is set to $0.001$ for both methods, and all other optimization protocols are kept the same as in experimental section. The LoRA configuration uses a rank of $8$ with a scaling factor ($\alpha$) of $32$, applies dropout at a rate of $0.05$, and targets both the query and value projection modules. As for IA3 method, adaptation is applied not only on query and value projection modules, but also for feed-forward layers.
B.2 Implementation of Transfer Learning Method
Algorithm 1 details the transfer learning method proposed for I2CL. In implementation, we empirically set $h=0.8$ and $\tau=0.5$.
@@ B.2 Implementation of Transfer Learning Method
Appendix C Additional Experiments and Analysis
C.1 Results under GPT2-XL, GPT-J, and Llama3-8B
Here, we further compare I2CL with baseline methods that do not need manual intervention or additional datasets on other two popular architectures. The results under architecture GPT-J and GPT2-XL are shown in Table 11 and Table 12, respectively. To highlight the generalization ability of I2CL, we further extend I2CL to latest Llama3-8B model and establish a comparison between zero-shot and few-shot learning in Table 13.
Table 11: Comparison between I2CL and baseline methods on GPT-J-6B.
@@ C.1 Results under GPT2-XL, GPT-J, and Llama3-8B
Table 12: Comparison between I2CL and baseline methods on GPT2-XL. AGnews and DBPedia are not evaluated due to the limitation of GPT2-XL’s context window size.
@@
Table 13: Performance comparison between Zero-shot, ICL, and I2CL on Llama3-8B.
Name SST-2 SST-5 TREC AGNews Subj HateSpeech18 DBPedia EmoC MR Avg. Zero-shot $55.60$ $31.40$ $66.40$ $73.60$ $51.00$ $50.40$ $56.20$ $41.00$ $55.60$ $53.47$ ICL $93.00\pm{0.62}$ $39.4\pm{3.21}$ $78.2\pm{5.94}$ $84.8\pm{1.30}$ $62.47\pm{12.44}$ $64.93\pm{2.65}$ $82.33\pm{3.64}$ $52.20\pm{3.92}$ $92.73\pm{0.50}$ $81.26$ I2CL $91.40\pm{2.26}$ $32.60\pm{1.25}$ $80.27\pm{2.58}$ $82.56\pm{1.57}$ $64.67\pm{3.52}$ $75.80\pm{1.02}$ $85.00\pm{0.33}$ $52.92\pm{5.28}$ $84.27\pm{2.07}$ $81.18$
C.2 Analysis on Synthetic Dataset
To underscore the generality of the proposed I2CL, we crafted a synthetic dataset that containing no semantic and formatting priors. Concretely, we generated three types of data containing “random strings”, “random special characters”, and “random numerical values”, labeled as “A” “B” and “C”, respectively. We then evaluate the performance of zero-shot, few-shot and I2CL on this synthetic dataset following the same setting as for main Table 1. As shown in Table 14, I2CL outperforms few-shot learning (i.e., ICL) by a large margin, highlighting the effectiveness and generality of I2CL, and we attribute this improvement to the proposed noisy self-calibration and vector injection methods which directly intervene at residual streams. Similar to the empirical observation in Sec. 3.3, the calibrated linear coefficients are also generalizable, and they can be directly used under unseen demonstration examples without additional calibration (see last column in Table 14).
Table 14: Evaluation of zero-shot, few-shot and I2CL on the synthetic dataset.
Task Zero-shot Few-shot (ICL) I2CL I2CL (unseen demo.) Synthetic data $32.6$ $66.20$$\pm$0.73 86.48$\pm$4.51 $86.36$$\pm$5.40
C.3 Apply I2CL over ICL
As I2CL and ICL are not mutually exclusive, an interesting empirical exploration involves applying I2CL on top of ICL. Specifically, we retain the demonstration tokens in the context throughout the noisy self-calibration and inference stages. As shown in Table 15, the combination of I2CL with ICL significantly improves performance on certain tasks, surpassing ICL by a substantial margin. Moreover, for tasks where I2CL originally underperforms compared to ICL, the inclusion of demonstration tokens helps bridge the gap. These empirical observations highlight the versatility of I2CL. On one hand, I2CL can serve as a substitute for ICL to reduce inference costs; on the other hand, when computational and memory constraints are less critical, I2CL can be applied in conjunction with ICL to further enhance ICL’s performance.
Table 15: Results of applying I2CL on top of ICL. Task AGNews is not applicable under 5-shot due to the limitation of GPU memory.
Name SST-2 SST-5 TREC Subj HateSpeech18 DBPedia EmoC MR Average ICL $\bf 94.44$$\pm$1.44 $41.72$$\pm$3.68 $77.32$$\pm$4.41 $52.56$$\pm$3.09 $70.24$$\pm$5.80 $\bf 96.64$$\pm$0.48 $\bf 75.48$$\pm$1.63 $\bf 93.24$$\pm$0.50 $75.21$ I2CL $87.68$$\pm$2.47 $39.12$$\pm$2.69 $78.56$$\pm$5.32 $73.84$$\pm$3.84 $69.88$$\pm$5.67 $90.16$$\pm$1.86 $63.72$$\pm$1.37 $87.68$$\pm$2.26 $73.83$ ICL+I2CL $93.52$$\pm$0.39 $\bf 44.53$$\pm$1.61 $\bf 83.56$$\pm$5.18 $\bf 89.84$$\pm$4.09 $\bf 81.36$$\pm$1.24 $95.87$$\pm$0.62 $73.84$$\pm$6.36 $93.16$$\pm$1.01 $\bf 81.96$
C.4 Additional Visualization
@@
Here, we present in Fig. 6 the calibrated coefficients for all tasks we used, illustrating variations and patterns across different configurations.
C.5 Visualization of In-task Context Vectors
As demonstrated in Fig. 4 (right), context vectors embed task semantics and context vectors from different classes are well separated. However, it is unclear whether context vector can carry label information within a more nuanced in-task perspective. To this end, we extract class-wise context vectors within each task and visualize in-task relations among different context vectors in Fig. 7. As shown in the figures, context vectors from different classes are well-separated in most tasks, indicating that context vectors will also carry label information. A notable exception comes from EmoC where vectors appear more mixed. We conjecture that besides label information, a substantial inherent parametric knowledge has also been hubbed by the context vector.
Appendix D Additional Discussion
D.1 Potential Design Space
Although the end residual stream contains a rich repository of information for the given token sequence, other token positions may also possess valuable attributes, as noted in recent studies (Hendel et al., 2023; Todd et al., 2024; Liu et al., 2024). Recent research also highlights the importance of formatting tokens during information propagation (Wang et al., 2023; Bai et al., 2024). We believe our approach could benefit from a more sophisticated design of demonstration vectorization. Additionally, while we currently use a single scalar to gauge the strength of aggregated multi-head attention, allowing finer granularity—such as separate strength scalars for different attention heads—might enhance our system’s performance, albeit at the cost of increased parameters.
D.2 Broader Impacts
I2CL inherits the same risks as standard In-context Learning. While I2CL primarily serves to enhance efficiency, its ability to adapt quickly to different data could potentially be used for creating misinformation or other harmful content at scale. I2CL’s reliance on existing data for demonstration examples could propagate existing biases if the data are not carefully curated. One additional risk comes from the generation of implicit vector representation for the demonstration examples, making the detection of the malicious contents even more challenging.
Figure 6: Calibrated coefficients for various datasets.
Figure 7: Visualization of in-task context vectors.